# Interpreting the Second-Order Effects of Neurons in CLIP

### v2

@@ Abstract
We interpret the function of individual neurons in CLIP by automatically describing them using text. Analyzing the direct effects (i.e. the flow from a neuron through the residual stream to the output) or the indirect effects (overall contribution) fails to capture the neurons’ function in CLIP. Therefore, we present the “second-order lens”, analyzing the effect flowing from a neuron through the later attention heads, directly to the output. We find that these effects are highly selective: for each neuron, the effect is significant for $<2\%$ of the images. Moreover, each effect can be approximated by a single direction in the text-image space of CLIP. We describe neurons by decomposing these directions into sparse sets of text representations. The sets reveal polysemantic behavior—each neuron corresponds to multiple, often unrelated, concepts (e.g. ships and cars). Exploiting this neuron polysemy, we mass-produce “semantic” adversarial examples by generating images with concepts spuriously correlated to the incorrect class. Additionally, we use the second-order effects for zero-shot segmentation and attribute discovery in images. Our results indicate that a scalable understanding of neurons can be used for model deception and for introducing new model capabilities. [^1] Project page and code: https://yossigandelsman.github.io/clip_neurons/
@@ 1 Introduction
Automated interpretability of the roles of components in neural networks enables the discovery of model limitations and interventions to overcome them. Recently, such a technique was applied for interpreting the attention heads in CLIP [13], a widely used class of image representation models [32]. However, this approach has only scratched the surface, failing to explain a major set of CLIP’s components—neurons. Here we will introduce a new interpretability lens for studying the neurons and use the gained understanding for segmentation, property discovery in images, and mass-production of semantic adversarial examples.
@@
Interpreting the neurons in CLIP is a harder task than interpreting the attention heads. First, there are more neurons than individual heads, which requires a more automated approach. Second, their direct effect on the output—the flow from the neuron, through the residual stream directly to the output—is negligible [13]. Third, most information is stored redundantly—many neurons encode the same concept, so just ablating a neuron (i.e. examining indirect effects) does not reveal much since other neurons make up for it.
@@
The limitations presented above mean that we can neither look at the direct effect nor the indirect effect to analyze a single neuron. To address this, we introduce a “second-order lens” for investigating the second-order effect of a neuron—its total contribution to the output, flowing via all the consecutive attention heads (see Fig. 1).
@@
We start by analyzing the empirical behavior of second-order effects of neurons. We find that these effects have high significance in the late layers. Additionally, each neuron is highly selective: its second-order effect is significant for only a small set (about 2%) of the images. Finally, this effect can be approximated by a single direction in the joint text-image representation space of CLIP (Sec. 3.3).
@@
As each direction that corresponds to a neuron lives in a joint representation space, it can be decomposed as a sparse sum of text representations that describes the neurons’ functionality (see Fig. 1). These text representations show that neurons are polysemantic [11]—each neuron corresponds to multiple semantic concepts. To verify that the neuron decompositions are meaningful, we show that these concepts correctly track which inputs activate a given neuron (Sec. 4).
@@
The polysemantic behavior of neurons allows us to find concepts that inadvertently overlap in the network, due to being represented by the same neuron. We use these spurious cues for mass production of “semantic” adversarial examples that will fool CLIP (see bottom of Fig. 1). We apply this technique to automatically produce adversarial images for a variety of classification tasks. Our qualitative and quantitative analysis shows that incorporating spuriously overlapping concepts in an image deceives CLIP with a significant success rate (Sec. 5.1).
@@
We present two additional applications – discovery of concepts in images, and zero-shot segmentation. For concept discovery, we interpret the concepts that CLIP associates with a given image by aggregating the text descriptions of the sparsely-activated neurons for that image. For segmentation, we can generate attribution heatmaps from the activation patterns of neurons. We do this by identifying class-relevant neurons with the second-order lens and averaging their heatmaps. This yields a strong zero-shot image segmenter that outperforms recent work [8, 13].
@@
In summary, we present an automated interpretability approach for CLIP’s neurons by modeling their second-order effects and spanning them with text descriptions. We use these descriptions to automatically understand neuron roles and apply this to three applications. This shows that a scalable understanding of internal mechanisms both uncovers errors and elicits new capabilities from models.
@@ 2 Related work
Contrastive vision-language models. Models like ALIGN [19], CLIP [32], and its variants [46, 22] produce image representations from pre-training on images and their captions. They demonstrated impressive zero-shot capabilities for various downstream tasks, including OCR, geo-localization, and classification [45]. These models’ representations are also used for segmentation [23], image generation [34, 36] and 3D understanding [20]. We aim to reveal the roles of neurons in such models.
Mechanistic interpretability of vision models. Mechanistic interpretability aims to reverse engineer the computation process in neural networks. In computer vision, this approach was applied to model individual network components [40] and to extract intermediate mechanisms like curve detectors [29], object segmenters [3, 2], high-frequency boundary detectors [37], and multimodal concepts detectors [15]. More closely to us, a few works made use of the intrinsic language-image space of CLIP to interpret the direct effect of attention heads and the output representation in CLIP with automatic text descriptions [13, 4]. We go beyond the output and direct effects of individual layers to interpret intermediate neurons in CLIP.
Neurons interpretability. The role of individual neurons (post-non-linearity single channel activations) is broadly studied in computer vision models [3, 2, 15] and language models [31, 14, 25]. [10, 17] demonstrate that neurons can learn universal mechanisms across different models in both domains. [11] show that neurons can be polysemantic (i.e. activated on multiple concepts) and exploit this property for generation of L2 adversarial examples. Some work seeks to extract neurons’ concepts by learning sparse dictionaries [7, 33]. Other methods use large language models to automatically describe neurons based on which examples they activate on [5, 28, 18, 41]. In contrast, we focus on the contribution of neurons to the output representation.
@@ 3 Second-order effects of neurons
We start by presenting the CLIP-ViT architecture. Then, we derive the second-order effect of neurons and present their benefits over first-order and the indirect effects. Finally, we empirically characterize the second-order effects, setting the stage for automatically interpreting them via text in Sec. 4.
@@ 3.1 CLIP-ViT Preliminaries
CLIP-ViT architecture. The CLIP-ViT image encoder consists of a Vision Transformer followed by a linear projection [^2] Throughout the paper, we ignore layer-normalization terms to simplify derivations. We address layers-normalization in detail in A.5.. The vision transformer (ViT) is applied to the input image $I\in\mathbb{R}^{H\times W\times 3}$ to obtain a $d^{\prime}$-dimensional representation $\mathsf{ViT}(I)$. Denoting the projection matrix by $P\in\mathbb{R}^{d\times d^{\prime}}$:
@@
The input $I$ to ViT is first split into $K$ non-overlapping image patches that are encoded into $K$ $d^{\prime}$-dimensional image tokens. An additional learned token, named the class token, is included and used later as the output token. As shown in Fig. 1, the tokens are processed simultaneously by applying $L$ alternating residual layers of multi-head self-attention (MSA) and MLP blocks.
@@
Figure 3: Mean-ablation of second order effects (ViT-B-32). We evaluate the performance on IN test set. Second-order effects concentrate in late layers, significant for only a part of the images, and can be approximated by one direction in the output space.
@@ 3.2 Analyzing the neuron effects on the output
Individual neurons have different types of contributions to the output—the first-order (direct) effects, second-order effects, and (higher-order) indirect effects. We introduce them and explain the limitations of the direct and indirect effects before continuing to characterize the second-order effects in Sec. 3.3.
First-order effects (logit lens [27]). The first-order effect is the direct contribution of a component to the residual stream, multiplied by the projection layer (see blue flow in Fig. 3). For an individual neuron $n$ in layer $l$, let $p_{i}^{l,n}(I)\in\mathbb{R}$ denote its post-GELU activation on the $i$-th token of the input image $I$. Then the contribution $e_{i}^{l,n}$ of the $n$-th neuron to the $i$-th token in the residual stream is:
@@
As observed by Gandelsman et al. [13], the first-order effects of MLP layers are close to constants in CLIP and most of the first-order contributions are from the late attention layers. We therefore focus on the second-order effects: the flow of information from the neurons through the attention layers.
Second-order effects. The contribution $e^{l,n}_{i}$ to the residual stream directly affects the input to later layers. We focus on the flow of $e^{l,n}_{i}$ through subsequent MSAs and then to the output (pink flow in Fig. 3). We call this interpretability lens the “second-order lens”, in analogy to the “logit lens”.
Following [12], the output of an MSA layer $\mathsf{MSA}^{l}$ that corresponds to the class token is a weighted sum of its $K+1$ input tokens $[z_{0},...,z_{K}]$:
@@
To obtain the second-order effect of a neuron $n$ at layer $l$, $\phi_{n}^{l}(I)$, we compute the additive contribution of the neuron through all the later MSAs and project it to the output space via $P$. Plugging in Eq. 3 as the contribution to $z_{i}$ in Eq. 4 and summing over layers, the second order effect of neuron $n$ is then:
@@
Indirect effects. An alternative approach is to analyze the indirect effect of a neuron by measuring the change in output representation when intervening on a neuron’s output. Specifically, the intervention is done by replacing the activation $p_{i}^{l,n}$ of the neuron for each token with a pre-computed per-token mean. However, as was shown by McGrath et al. [24], models often learn “self-repair” mechanisms that can obscure the individual roles of neurons. We illustrate these issues in the next section.
@@
Figure 4: Comparison to indirect effect. We compare the second-order effects and the indirect effects by mean-ablating layer 9 in ViT-B-32 on ImageNet test set.
@@ 3.3 Characterizing the second-order effects
We analyze the empirical behavior of the second-order effects of neurons $\phi^{l}_{n}$ derived in the previous section. We find that only neurons from the late MLP layers have a significant second-order effect and that each individual neuron has a significant effect for less than 2% of the images. Finally, we show that $\phi^{l}_{n}$ can be approximated by one linear direction in the output space. These findings will help motivate our algorithm for describing output spaces of neurons with text in Sec. 4.
Experimental setting. To evaluate the second-order effects and their contributions to the output representation, we measure the downstream performance on the ImageNet classification task [9] after ablating these effects for each neuron. Specifically, we apply mean-ablation [26], replacing the additive contributions of individual $\phi_{n}^{l}(I)$’s to the representation with the mean computed across a dataset $D$. In our experiments, we mean-ablate all the neurons in a layer simultaneously and evaluate the downstream classification performance before and after ablation. Components with larger effects should result in larger accuracy drops.
We take $D$ to be $\sim\!5000$ images from the ImageNet (IN) training set. We report zero-shot classification accuracy on the IN test set. Our model is OpenAI’s ViT-B-32 CLIP, which has 12 layers (additional results for ViT-L-14 are in App. A.1)
Second-order effects concentrate in moderately late layers. We evaluate the contributions of all the $\phi^{l}_{n}$ across different layers and observe that the neurons with the most significant second-order effects appear relatively late in the model. The results for different layers in ViT-B-32 CLIP model are presented in Fig. 3 (“w/o all neurons”). As shown, mean-ablating layers 8-10 leads to the largest drop in performance. These layers appear right before the MSA layers with the most significant direct effects, as shown in [13] (layers 9-11; see App. A.2).
The second-order effect is sparse. We find that the second-order effect of each individual neuron is significant only for less than 2% of the images across the test set. We repeat the same experiment as before, but this time we only mean-ablate $\phi^{l}_{n}(I)$ for a subset of images, while keeping the original effects for other images. We find that for most of the images, except the subset of images in which $\phi^{l}_{n}(I)$ has a large norm, we can mean-ablate $\phi^{l}_{n}(I)$ without changing the accuracy significantly, as shown in Fig. 3 (“w/o small norm”). Differently, mean-ablating the contributions for the 100 images with the largest $\phi^{l}_{n}(I)$ norms results in a significant drop in performance (“w/o large norm”).
@@
To verify that this approximation recovers $\phi^{l}_{n}(I)$ we replace each $\phi^{l}_{n}(I)$ for each neuron and image in the test set with the approximation. We then evaluate the downstream classification performance. As shown in Fig. 3 (“reconstruction from PC #1”), for each layer $l$, this replacement results in a negligible drop in performance from the baseline, that uses the full representation.
Comparison to indirect effect. We compare the second-order effect to the indirect effect and present the variance explained by the first principle component for each of them and the drop in performance when simultaneously mean-ablating all the effects from one layer. As shown in Tab. 5, Mean-ablating the indirect effects results in a smaller drop in performance due to self-repair behavior. Moreover, the first principle component explains significantly less of the variance in the indirect effect, than in the second-order effect. This demonstrates two advantages of the second-order effects—uncovering neuron functionality that is obfuscated by self-repair, and one-dimensional behavior that can be easily modeled and decomposed, as we will show in the next section.

### v3

@@ Abstract
We interpret the function of individual neurons in CLIP by automatically describing them using text. Analyzing the direct effects (i.e. the flow from a neuron through the residual stream to the output) or the indirect effects (overall contribution) fails to capture the neurons’ function in CLIP. Therefore, we present the “second-order lens”, analyzing the effect flowing from a neuron through the later attention heads, directly to the output. We find that these effects are highly selective: for each neuron, the effect is significant for $<2\%$ of the images. Moreover, each effect can be approximated by a single direction in the text-image space of CLIP. We describe neurons by decomposing these directions into sparse sets of text representations. The sets reveal polysemantic behavior—each neuron corresponds to multiple, often unrelated, concepts (e.g. ships and cars). Exploiting this neuron polysemy, we mass-produce “semantic” adversarial examples by generating images with concepts spuriously correlated to the incorrect class. Additionally, we use the second-order effects for zero-shot segmentation, outperforming previous methods. Our results indicate that an automated interpretation of neurons can be used for model deception and for introducing new model capabilities [^1] Project page: https://yossigandelsman.github.io/clip_neurons/..
@@ 1 Introduction
Automated interpretability of the roles of components in neural networks enables the discovery of model limitations and interventions to overcome them. Recently, such a technique was applied for interpreting the attention heads in CLIP (Gandelsman et al., 2024), a widely used class of image representation models (Radford et al., 2021). However, this approach has only scratched the surface, failing to explain a major set of CLIP’s components—neurons. Here we will introduce a new interpretability lens for studying the neurons and use the gained understanding for zero-shot segmentation and mass-production of semantic adversarial examples.
@@
Interpreting the neurons in CLIP is a harder task than interpreting the attention heads. First, there are more neurons than individual heads, which requires a more automated approach. Second, their direct effect on the output—the flow from the neuron, through the residual stream directly to the output—is negligible (Gandelsman et al., 2024). Third, most information is stored redundantly—many neurons encode the same concept, so just ablating a neuron (i.e. examining indirect effects) does not reveal much since other neurons make up for it.
@@
The limitations presented above mean that we can neither look at the direct effect nor the indirect effect to analyze a single neuron. To address this, we introduce a “second-order lens” for investigating the second-order effect of a neuron—its total contribution to the output, flowing via all the consecutive attention heads (see Figure 1).
@@
We start by analyzing the empirical behavior of second-order effects of neurons. We find that these effects have high significance in the late layers. Additionally, each neuron is highly selective: its second-order effect is significant for only a small set (about 2%) of the images. Finally, this effect can be approximated by a single direction in the joint text-image representation space of CLIP (Section 3.3).
@@
As each direction that corresponds to a neuron lives in a joint representation space, it can be decomposed as a sparse sum of text representations that describes the neurons’ functionality (see Figure 1). These text representations show that neurons are polysemantic (Elhage et al., 2022)—each neuron corresponds to multiple semantic concepts. To verify that the neuron decompositions are meaningful, we show that these concepts correctly track which inputs activate a given neuron (Section 4).
@@
The polysemantic behavior of neurons allows us to find concepts that inadvertently overlap in the network, due to being represented by the same neuron. We use these spurious cues for mass production of “semantic” adversarial examples that will fool CLIP (see bottom of Figure 1). We apply this technique to automatically produce adversarial images for a variety of classification tasks. Our qualitative and quantitative analysis shows that incorporating spuriously overlapping concepts in an image deceives CLIP with a significant success rate (Section 5.1).
@@
The text representations that describes the neurons’ functionality enable an additional application—zero-shot segmentation. Mining for text representations of class names, we can identify class-relevant neurons with the second-order lens. Averaging the activation patterns of such neurons, we generate attribution heatmaps. Binarizing them yields a strong zero-shot image segmenter that outperforms recent work (Chefer et al., 2021; Gandelsman et al., 2024).
@@
In summary, we present an automated interpretability approach for CLIP’s neurons by modeling their second-order effects and spanning them with text descriptions. We use these descriptions to automatically understand neuron roles and apply this to two applications. This shows that a scalable understanding of internal mechanisms both uncovers errors and elicits new capabilities from models.
@@ 2 Related work
Contrastive vision-language models. Models like ALIGN (Jia et al., 2021), CLIP (Radford et al., 2021), and its variants (Zhai et al., 2023; Li et al., 2023) produce image representations from pre-training on images and their captions. They demonstrated impressive zero-shot capabilities for various downstream tasks, including OCR, geo-localization, and classification (Wortsman, 2023). These models’ representations are also used for segmentation (Lüddecke & Ecker, 2022), image generation (Ramesh et al., 2021; Rombach et al., 2022) and 3D understanding (Kerr et al., 2023). We aim to reveal the roles of neurons in such models.
Mechanistic interpretability of vision models. Mechanistic interpretability aims to reverse engineer the computation process in neural networks. In computer vision, this approach was applied to model individual network components (Shah et al., 2024) and to extract intermediate mechanisms like curve detectors (Olah et al., 2020), object segmenters (Bau et al., 2019; 2020), high-frequency boundary detectors (Schubert et al., 2021), and multimodal concepts detectors (Goh et al., 2021). More closely to us, a few works made use of the intrinsic language-image space of CLIP to interpret the direct effect of attention heads and the output representation in CLIP with automatic text descriptions (Gandelsman et al., 2024; Bhalla et al., 2024). We go beyond the output and direct effects of individual layers to interpret intermediate neurons in CLIP.
Neurons interpretability. The role of individual neurons (post-non-linearity single channel activations) is broadly studied in computer vision models (Bau et al., 2019; 2020; Goh et al., 2021) and language models (Radford et al., 2017; Geva et al., 2021; Meng et al., 2022). (Dravid et al., 2023; Gurnee et al., 2024) demonstrate that neurons can learn universal mechanisms across different models in both domains. Elhage et al. (2022) show that neurons can be polysemantic (i.e. activated on multiple concepts) and exploit this property for generation of L2 adversarial examples. Some work seeks to extract neurons’ concepts by learning sparse dictionaries (Bricken et al., 2023; Rajamanoharan et al., 2024). Other methods use large language models to automatically describe neurons based on which examples they activate on (Bills et al., 2023; Oikarinen & Weng, 2023; Hernandez et al., 2022; Shaham et al., 2024). In contrast, we focus on the contribution of neurons to the output representation.
@@ 3 Second-order effects of neurons
We start by presenting the CLIP-ViT architecture. Then, we derive the second-order effect of neurons and present their benefits over first-order and the indirect effects. Finally, we empirically characterize the second-order effects, setting the stage for automatically interpreting them via text in Section 4.
@@ 3.1 CLIP-VIT Preliminaries
CLIP-ViT architecture. The CLIP-ViT image encoder consists of a Vision Transformer followed by a linear projection [^2] Throughout the paper, we ignore layer-normalization terms to simplify derivations. We address layers-normalization in detail in Section A.6.. The vision transformer (ViT) is applied to the input image $I\in\mathbb{R}^{H\times W\times 3}$ to obtain a $d^{\prime}$-dimensional representation $\mathsf{ViT}(I)$. Denoting the projection matrix by $P\in\mathbb{R}^{d\times d^{\prime}}$:
@@
The input $I$ to ViT is first split into $K$ non-overlapping image patches that are encoded into $K$ $d^{\prime}$-dimensional image tokens. An additional learned token, named the class token, is included and used later as the output token. As shown in Figure 1, the tokens are processed simultaneously by applying $L$ alternating residual layers of multi-head self-attention (MSA) and MLP blocks.
@@
Figure 3: Mean-ablation of second order effects (ViT-B-32). We evaluate the performance on ImageNet validation set. Second-order effects concentrate in late layers, significant for only a part of the images, and can be approximated by one direction in the output space.
@@ 3.2 Analyzing the neuron effects on the output
Individual neurons have different types of contributions to the output—the first-order (direct) effects, second-order effects, and (higher-order) indirect effects. We introduce them and explain the limitations of the direct and indirect effects before continuing to characterize the second-order effects in Section 3.3.
First-order effects (logit lens (nostalgebraist, 2020)). The first-order effect is the direct contribution of a component to the residual stream, multiplied by the projection layer (see blue flow in Figure 3). For an individual neuron $n$ in layer $l$, let $p_{i}^{l,n}(I)\in\mathbb{R}$ denote its post-GELU activation on the $i$-th token of the input image $I$. Then the contribution $e_{i}^{l,n}$ of the $n$-th neuron to the $i$-th token in the residual stream is:
@@
As observed by Gandelsman et al. (2024), the first-order effects of MLP layers are close to constants in CLIP and most of the first-order contributions are from the late attention layers. We therefore focus on the second-order effects: the flow of information from the neurons through the attention layers.
Second-order effects. The contribution $e^{l,n}_{i}$ to the residual stream directly affects the input to later layers. We focus on the flow of $e^{l,n}_{i}$ through subsequent MSAs and then to the output (pink flow in Figure 3). We call this interpretability lens the “second-order lens”, in analogy to the “logit lens”.
Following Elhage et al. (2021), the output of an MSA layer $\mathsf{MSA}^{l}$ that corresponds to the class token is a weighted sum of its $K+1$ input tokens $[z_{0},...,z_{K}]$:
@@
To obtain the second-order effect of a neuron $n$ at layer $l$, $\phi_{n}^{l}(I)$, we compute the additive contribution of the neuron through all the later MSAs and project it to the output space via $P$. Plugging in Equation 3 as the contribution to $z_{i}$ in Equation 4 and summing over layers, the second order effect of neuron $n$ is then:
@@
Indirect effects. An alternative approach is to analyze the indirect effect of a neuron by measuring the change in output representation when intervening on a neuron’s output. Specifically, the intervention is done by replacing the activation $p_{i}^{l,n}$ of the neuron for each token with a pre-computed per-token mean. However, as was shown by McGrath et al. (2023), models often learn “self-repair” mechanisms that can obscure the individual roles of neurons. We illustrate these issues in the next section.
@@
Figure 4: Comparison to indirect effect. We compare the second-order effects and the indirect effects by mean-ablating layer 9 in ViT-B-32 on ImageNet validation set.
@@ 3.3 Characterizing the second-order effects
We analyze the empirical behavior of the second-order effects of neurons $\phi^{l}_{n}$ derived in the previous section. We find that only neurons from the late MLP layers have a significant second-order effect and that each individual neuron has a significant effect for less than 2% of the images. Finally, we show that $\phi^{l}_{n}$ can be approximated by one linear direction in the output space. These findings will help motivate our algorithm for describing output spaces of neurons with text in Section 4.
Experimental setting. To evaluate the second-order effects and their contributions to the output representation, we measure the downstream performance on the ImageNet classification task (Deng et al., 2009) after ablating these effects for each neuron. Specifically, we apply mean-ablation (Nanda et al., 2023), replacing the additive contributions of individual $\phi_{n}^{l}(I)$’s to the representation with the mean computed across a dataset $D$. In our experiments, we mean-ablate all the neurons in a layer simultaneously and evaluate the downstream classification performance before and after ablation. Components with larger effects should result in larger accuracy drops.
We take $D$ to be $\sim\!5000$ images from the ImageNet training set. We report zero-shot classification accuracy on the ImageNet validation set. Our model is OpenAI’s ViT-B-32 CLIP, which has 12 layers. We present additional results for ViT-L-14 and for ImageNet-R (Hendrycks et al., 2021) in Section A.1 and Figure 11.
Second-order effects concentrate in moderately late layers. We evaluate the contributions of all the $\phi^{l}_{n}$ across different layers and observe that the neurons with the most significant second-order effects appear relatively late in the model. The results for different layers in ViT-B-32 CLIP model are presented in Figure 3 (“w/o all neurons”). As shown, mean-ablating layers 8-10 leads to the largest drop in performance. These layers appear right before the MSA layers with the most significant direct effects, as shown in Gandelsman et al. (2024) (layers 9-11; see Section A.2). The same trend is preserved for a larger model size as well (see Section A.1).
The second-order effect is sparse. We find that the second-order effect of each individual neuron is significant only for less than 2% of the images across the validation set. We repeat the same experiment as before, but this time we only mean-ablate $\phi^{l}_{n}(I)$ for a subset of images, while keeping the original effects for other images. For most of the images, except the subset of images in which $\phi^{l}_{n}(I)$ has a large norm, we can mean-ablate $\phi^{l}_{n}(I)$ without changing the accuracy significantly, as shown in Figure 3 (“w/o small norm”). Differently, mean-ablating the contributions for the 100 images with the largest $\phi^{l}_{n}(I)$ norms results in a significant drop in performance (“w/o large norm”). The same trend is shown for images from ImageNet-R in Figure 11.
@@
To verify that this approximation recovers $\phi^{l}_{n}(I)$ we replace each $\phi^{l}_{n}(I)$ for each neuron and image in the validation set with the approximation. We then evaluate the downstream classification performance. As shown in Figure 3 (“reconstruction from PC #1”), for each layer $l$, this replacement results in a negligible drop in performance from the baseline, that uses the full representation. The same behavior is observed for ViT-L model and for a different initial set of images in the Appendix.
Comparison to indirect effect. We compare the second-order effect to the indirect effect and present the variance explained by the first principle component for each of them and the drop in performance when simultaneously mean-ablating all the effects from one layer. As shown in Figure 5, Mean-ablating the indirect effects results in a smaller drop in performance due to self-repair behavior. Moreover, the first principle component explains significantly less of the variance in the indirect effect, than in the second-order effect. This demonstrates two advantages of the second-order effects—uncovering neuron functionality that is obfuscated by self-repair, and one-dimensional behavior that can be easily modeled and decomposed, as we will show in the next section.